OpenStudy (anonymous):

whats the sqrt of 24 over 18 plz help

OpenStudy (tkhunny):

What have you tried? \(\sqrt{24} = \sqrt{4\cdot 6} = ??\)

OpenStudy (anonymous):

sqrt[24/18] = sqrt[4/3] = 2/sqrt[3] = 2/sqrt[3] * sqrt[3] / sqrt[3] = (2/3)*sqrt[3]

OpenStudy (anonymous):

or simplfy as sqrt(24/18) = sqrt(4/3) = 2/sqrt(3)

OpenStudy (anonymous):

u can't leave the denominator with sqrt so u need to multiply 2/sqrt[3] * sqrt[3] / sqrt[3] to get (2/3)*sqrt[3]

OpenStudy (tkhunny):

You should specify more clearly. Where are we starting? \(\dfrac{\sqrt{24}}{18}\;or\;\sqrt{\dfrac{24}{18}}\)